On Friday , Garcia traveled 363 miles from tiptonville to oak ridge in 7.5 hours . if he travels at the same rate on Saturday , how far will he travel in 5 hours
Answer by TimothyLamb(4379)

s = d/t
s = 363/7.5
---
d = st
d = (363/7.5)*5
d = 242 miles
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php